OpenStudy (anonymous):

Carla runs 4 miles every day. she jogs from home to the school track witch is 3/4 mile away. she then runs laps around the 1/4 mile track Carla then jogs home . how many laps does she run at the school?

OpenStudy (anonymous):

so she runs 13

OpenStudy (anonymous):

thank you

OpenStudy (pokemon23):

13 for sure just follow moneybird steps in solving it

hero (hero):

Seems like only 10 laps to me

OpenStudy (pokemon23):

Hero?

OpenStudy (pokemon23):

Hero is going to school us

OpenStudy (anonymous):

6/4 + 1/4 x = 4

hero (hero):

4 - 1.50 = 2.50 2.50 = 8 + 2 quarter milers

OpenStudy (anonymous):

6 + x = 16 x = 10

hero (hero):

Wink

OpenStudy (anonymous):

Carla is obviously trying to loose weight

OpenStudy (anonymous):

4-1.5=2.5/.25=10 [4miles-1.5 miles round trip, leaves 2 1/2 miles\1/4 mile=10 laps]

hero (hero):

Somehow, I missed getting a medal :/

hero (hero):

Thanks :D

OpenStudy (anonymous):

Np :D